OpenStudy (anonymous):

Please help, will give medal and follow if this can be solved properly! I am looking for the answer to part x.

OpenStudy (anonymous):

In the figure below, angle y and angle x form vertical angles. Angle x forms a straight line with the 50° angle and the 40° angle. Write and solve an equation

OpenStudy (anonymous):



OpenStudy (anonymous):

So overall there are 360 degrees in this diagram so 360=x+y+180. the 180 comes from the 50+40 degrees on the left plus the 90 degrees to the right which is adjacent to y and x.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so this is the problem: 360 = x + y +180 50 + 40 + 90

OpenStudy (anonymous):

Is 90 the answer?

OpenStudy (anonymous):

Yes 180= x+y and x=y= 40+50

OpenStudy (anonymous):

Okay, Thank you!